solve 3y^2+4y-2<= 0
Answer by stanbon(75887)

solve 3y^2+4y-2<= 0
-------------------------
Solve the EQUALITY:
y = [-4 +- sqrt(16 - 4*3*-2)]/6
y = [-4 +- sqrt(40)]/6
y = [-4 +- 2sqrt(10)]/6
y = (-2/3)+(1/3)sqrt(10) or y = (-2/3)-(1/3)sqrt(10)
Approximate solutions: y = 0.3874 or y = -1.721
=====================================================
Solve the INEQUALITY:
Draw a number line. Plot the two equality solutions.
Check a value in each of the three intervals formed by those solutions
to see where the solution for the inequality lie.
--------------------
3y^2+4y-2<= 0
Check y = -10 ; 3(-10)^2+4(-10)-2 < 0 ; false
Check y = 0 ; -2 < 0 ; true so solutions in (-1.721,0.3874)
check y = 10 ; 3(10)^2-4(10)-2 < 0 ; false
-------------------------------------------------
Final solution; (-2/3)-(1/3)sqrt(10) <= y <= (-2/3)+(1/3)sqrt(10)
============================================================================
Cheers,
Stan H.